OpenStudy (anonymous):

The temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later.

OpenStudy (anonymous):

both eqautions equaled 53 when i plug in 6?is that correct?

OpenStudy (anonymous):

Firstly both the equation are one and the same thing.. You can write the equation as: \[y = -2x + 65\] Or you can write it as: \[y = 65 -2x\] Both are same..

OpenStudy (anonymous):

Yes for finding temperature 6 hours later you will put x = 6; \[y = 65 - 2(6) = \] Yes your answer is correct..

OpenStudy (anonymous):

y = 53 Your answer is correct..

OpenStudy (anonymous):

thank you so much!i needed to make sure :)

OpenStudy (anonymous):

Welcome dear..